Write the following in scientific notation: 1.4x10^-2/5.6x10^-6
Answer by edjones(8007)

1.4x10^-2/5.6x10^-6
=5.6*10^6 / 1.4*10^2
=5.6*10^4/1.4
=4*10^4
.
Ed